after The decrease in the price of peas to 20N,assume that John is just as well of as he was in (a) if he has an income of zmw7.6.However,with that income and the new price of peas he would have consumed 20kg of beans.find the quantity of peas he would have consumed in this case.
1
Expert's answer
2020-09-28T07:50:53-0400

After decrease of peas to 20 N per kg,


New Income, we use the new income =ZMW 7.6 to evaluate the change in quantity of peas


We get the new total income for peas only =760-(20×20)=360


Quantity ="\frac{360}{20}" =18Kg


Quantity of peas =18Kg